# Mean field model of contagion processes in urban traffic networks

### v1

@@ 3 Comparisons with empirical data
Our contagion network traffic flow model has been tested on representative instances of the publicly available UTD19 database [4]. This database provides fundamental traffic variables including flow and density, from 41 cities worldwide. The data is collected from stationary traffic sensors, most of them inductive loop detectors, which report traffic flow and the fraction of time during an observation period in which a detector is occupied as a proxy for the vehicle density.
@@
The network contagion model predicts a SIR dynamics governing the vehicle flow over $24$ hour periods. The upper and lower panels of the Figure 2, show the fitting of the SIR density dynamics from a single detector over a particular street, in the Aussersihl district of Zurich, Switzerland. This instance of the UTD19 database has been showcased in a number of works and has good quality sensor data [4]. Fitting of the SIR rate parameters is done by minimizing the total squared error between the observed densities and those predicted by the SIR dynamics integrated by first order discretization through the Euler method. Because the street density data is a proxy, additional parameters for a linear transformation that resolves the particular sensor calibration are introduced. The resulting optimization problem has two levels. Firstly the rate parameters are estimated by solving
@@
The calibration parameter $10<s_{0}<$ 1 estimates the fraction of the overall transport covered by the $n$-th link in the mean field limit. The parameters $c_{1}$ and $c_{2}$ are mere calibration constants for the rates, $\beta_{n}\to c_{1}\beta_{n}$, $\gamma_{n}\to c_{2}\gamma_{n}$.
@@
Figure 3: Predicted (red triangles) and observed (black circles) transport in 22 detectors from the Aussersihl district in Zurich, Switzerland, at Wednesday 28th October of 2015. From the estimated transmission rate parameters, the mean field crossroad contagion network traffic flow model predicts an average road velocity over the covered region of $30.92$ km/hr and an average venting rate of $6.54$ km/hr. The arbitraily chosen detectors from the UTD19 database are, K14D15, K14D13, K14D14, K17D13, K17D14, K17D15, K17D12, K17D11, K13D14, K13D13, K13D12, K14D12, K11D19, K11D11, K12D16, K133D11, K133D12, K133D13, K11D12, K11D14, K12D14, K12D17.
@@ 4 Discussion
Despite being an idealization, the contagion framework to traffic appears to capture features that are essential to characterize urban transport at different spatial and temporal scales. The formalism integrates granular with continuous aspects of vehicle flow in conjunction with basic temporal constraints, leading to a local mean field contagion type model. Our empirical comparisons indicate that single detector data aggregates information from all of the links directly connected to the detector’s road. The macroscopic mean field equation (16), predicts that aggregating links in a given macroscopic area of the roads network should converge to macroscopic rate parameters that reflect the vehicle flow in the corresponding network’s section. This appears to be consistent with the limited empirical study reported in the Figure 3. A more thorough experimental study is currently under development by the authors of the present article.
5 Conclusions and future work
In this work it has been introduced a mean field model for vehicle flow in traffic networks that takes into account granular features of the flow at the link’s junctions. This is accomplished by a SIR-type contagion dynamics that aggregates the vehicles sent to a given link. The link level rate parameters are determined by the total transport that traverse the link in a given time interval. We have at this point provided comparisons with empirical data that support our model at individual links and in groups of not necessarily directly connected links inside a macroscopic zone of the network. Clearly a next important step is to apply our approach to large macroscopic traffic networks datasets, by carefully taking into account the network topology. Our equation (18) indicates that the free flow optimal state of a macroscopic traffic network occurs in a critical density given in terms of aggregated rate parameters $\beta$, $\gamma$ and a normalization parameter associated with the total transport in the given time period, $s_{0}$.
Generalizations to the mean field setup that consider network’s spatiotemporal structures in greater detail, can in principle be explored by taking advantage of recent developments in the study of contagion models on networks with complex topologies, like for instance those presented in [10].
@@ 6 Acknowledgments
The authors acknowledge to the creators and maintainers of the UTD19 database. The UTD19 database is hosted in the site utd19.ethz.ch

### v3

@@ 3.2 Comparisons with traffic sensors data
Our contagion network traffic flow model has also been tested on representative instances of the publicly available UTD19 database [loder]. This database provides fundamental traffic variables including flow and density, from 41 cities worldwide. The data is collected from stationary traffic sensors, most of them inductive loop detectors, which report traffic flow and the fraction of time during an observation period in which a detector is occupied as a proxy for the vehicle density.
@@
The upper and lower panels of the Figure 5, show the fitting of the SIR density dynamics from a single detector over a particular street, in the Aussersihl district of Zurich, Switzerland, over a $24$ hour period. This instance of the UTD19 database has been showcased in a number of works and has good quality sensor data [loder]. Fitting of the SIR rate parameters is done by minimizing the total squared error between the observed densities and those predicted by the SIR dynamics integrated by first order discretization through the Euler method. Because the street density data is a proxy, additional parameters for a linear transformation that resolves the particular sensor calibration are introduced. The resulting optimization problem has two levels. Firstly the rate parameters are estimated by solving
@@
The calibration parameter $0<s_{0}<1$ estimates the fraction of the overall transport covered by the $n$-th link in the mean field limit. The parameters $c_{1}$ and $c_{2}$ are mere calibration constants for the rates, $\beta_{n}\to c_{1}\beta_{n}$, $\gamma_{n}\to c_{2}\gamma_{n}$.
@@
Figure 6: Predicted (red triangles) and observed (black circles) transport in 22 detectors from the Aussersihl district in Zurich, Switzerland, at Wednesday 28th October of 2015. From the estimated transmission rate parameters, the mean field crossroad contagion network traffic flow model predicts an average inflow over the covered region of $30.92$ veh/hr and an average venting rate of $6.54$ veh/hr. The arbitraily chosen detectors from the UTD19 database are, K14D15, K14D13, K14D14, K17D13, K17D14, K17D15, K17D12, K17D11, K13D14, K13D13, K13D12, K14D12, K11D19, K11D11, K12D16, K133D11, K133D12, K133D13, K11D12, K11D14, K12D14, K12D17.
@@ 4 Discussion
Despite being an idealization, the contagion framework to traffic appears to capture features that are essential to characterize urban transport at different spatial and temporal scales. The formalism integrates granular with continuous aspects of vehicle flow in conjunction with basic temporal constraints, leading to a local mean field contagion type model. Our empirical comparisons indicate that single detector data aggregates information from all of the links directly connected to the detector’s road. In fact, the experiment presented in Subsection 3.1 indicates that it’s possible to estimate macroscopic traffic network parameters, like total transport or optimal vehicle transit rates over a region, from purely local flow measurements at an intersection.
The macroscopic mean field equation (8), predicts that aggregating links in a given macroscopic area of the roads network should converge to macroscopic rate parameters that reflect the vehicle flow in the corresponding network’s section. This appears to be consistent with the empirical study reported in the Figure 6.
Clearly a next important step in the development of the proposed mean field theory, is to apply our approach to large macroscopic traffic networks datasets, by carefully taking into account the network topology. Our equation (10) indicates that the free flow optimal state of a macroscopic traffic network occurs in a critical density given in terms of aggregated rate parameters $\beta$, $\gamma$ and a normalization parameter associated with the total transport in the given time period, $s_{0}$. This may lead to “immunization” traffic control strategies, as explained in Subsection 2.1.
Generalizations to the mean field setup that consider network’s spatiotemporal structures in greater detail, can in principle be explored by taking advantage of recent developments in the study of contagion models on networks with complex topologies, like for instance those presented in [palafox].
@@ 5 Acknowledgments
The authors are grateful with Dr. Jaime Burgos García and Dr. Simón Rodríguez Rodríguez for the very useful discussions with them and particularly to Dr. Burgos García for having kindly shared with us the experimental data from the Municipal Institute of Sustainable Urban Mobility of Saltillo analyzed by us in Subsection 3.1.
The authors also acknowledge to the creators and maintainers of the UTD19 database, from which the traffic sensors data used in Subsection 3.2 has been taken. The UTD19 database is hosted in the site utd19.ethz.ch